# Robust Five-Class and binary Diabetic Retinopathy Classification Using Transfer Learning and Data Augmentation

### v1

@@ Abstract
A robust deep learning framework is proposed for both binary and five-class diabetic retinopathy (DR) classification.
Class-balanced data augmentation strategy significantly improves model performance and addresses dataset imbalance.
Extensive evaluation of pretrained CNNs (ResNet, EfficientNet) identifies optimal architectures for DR classification.
Achieves state-of-the-art binary classification accuracy of 98.9% with an AUC of 99.4%.
Attains competitive performance in five-class DR severity grading with 84.6% accuracy and 94.1% AUC.
@@ 3.1 Data Acquisition and Preprocessing
Images were obtained from the APTOS 2019 dataset and split into training, validation, and test subsets. All images were converted to RGB format and resized to $224\times 224$ pixels to maintain consistency. The images were loaded into memory using the Python Imaging Library (PIL), and pixel values were normalized using ImageNet statistics. The datasets were combined and split with an 85:15 ratio using stratified sampling to preserve class distributions. The number of samples per class in the training and test sets, both before and after data augmentation, is presented in Table 1 and Table 2.
@@ 3.2 Data Augmentation
  • RandomResizedCrop to $224\times 224$ with scale in $[0.7,1.0]$
@@ 4.1 Dataset
APTOS 2019 dataset was used for a Kaggle competition on DR diagnosis [16]. The images have varying resolutions, ranging from $474\times 358$ to $3388\times 2588$. APTOS stands for Asia Pacific Tele-Ophthalmology Society, and the dataset was provided by Aravind Eye Hospital in India. In this dataset, fundus images are graded manually on a scale of 0 to 4 (0: no DR; 1: mild; 2: moderate; 3: severe; and 4: proliferative DR) to indicate different severity levels. The number of images in these classes are respectively 1805, 370, 999, 193, and 295. In the binary setting, class 0 is defined as the normal group, and the remaining classes (1-4) are defines as DR group which gives a split 1805:1857. The total number of training and test samples in the dataset were 3662 and 1928 respectively. However, the labels for the test samples were not released after the competition, so like other references, we used the available 3662 fundus images with labels. To ensure fair comparison with existing deep learning approaches, we employ a standard majority-based train–test split protocol. The class-wise distribution of samples in the training and test sets—before and after augmentation—is summarized in Table 1 for the binary classification setup and in Table 2 for the five-class (multiclass) setting.
@@ 4.2 Hyperparameters
All models were trained using the Adam optimizer with a fixed learning rate of $1\times 10^{-4}$ and a batch size of 32. Input images were resized to $224\times 224$ pixels and normalized using ImageNet statistics. Cross-entropy loss was used for both binary and multiclass classification tasks. Training was performed for a maximum of 500 epochs, with early stopping applied if the macro F1-score on the validation set did not improve for 50 consecutive epochs. During augmentation, each class in the training set was expanded to 20,000 samples to address class imbalance.
@@ 9 Future Work
MANB reviewed and edited the manuscript. FA downloaded the data, conceptualized the study, prepared the code, performed the data analysis and wrote the manuscript. Both authors reviewed and approved the final version of the manuscript.
@@
This work was supported by the Institutional Development Award (IDeA) from the National Institutes of General Medical Sciences of the NIH under grant number P20GM121307 to MANB.

### v2

# Addressing High Class Imbalance in Multi-Class Diabetic Retinopathy Severity Grading with Augmentation and Transfer Learning

@@ Abstract
A deep learning framework is developed for diabetic retinopathy (DR) severity classification, addressing both binary and five-class tasks.
A class-balanced data augmentation strategy is introduced, generating 20,000 synthetic samples per class to mitigate data imbalance.
Transfer learning and fine-tuning of pretrained CNNs (EfficientNet, ResNet) are extensively evaluated to identify optimal architectures for DR classification.
The proposed model achieves state-of-the-art binary classification performance with 98.9% accuracy and 99.4% AUC.
Competitive results are obtained for five-class DR grading, reaching 84.6% accuracy and 94.1% AUC.
@@ 3.1 Data Acquisition and Preprocessing
Images were obtained from the APTOS 2019 dataset and split into training, validation, and test subsets. All images were converted to RGB format and resized to $\times 224$ 224 pixels to maintain consistency. The images were loaded into memory using the Python Imaging Library (PIL), and pixel values were normalized using ImageNet statistics. The datasets were combined and split with an 85:15 ratio using stratified sampling to preserve class distributions. The number of samples per class in the training and test sets, both before and after data augmentation, is presented in Table 1 and Table 2.
@@ 3.2 Data Augmentation
  • RandomResizedCrop to $\times 224$ 224 with scale in [0.7,1.0][0.7,1.0][ 0.7 , 1.0 ]
@@ 4.1 Dataset
APTOS 2019 dataset was used for a Kaggle competition on DR diagnosis [16]. The images have varying resolutions, ranging from 474×358474$\times$ 358474 × 358 to 3388×25883388$\times$ 25883388 × 2588. APTOS stands for Asia Pacific Tele-Ophthalmology Society, and the dataset was provided by Aravind Eye Hospital in India. In this dataset, fundus images are graded manually on a scale of 0 to 4 (0: no DR; 1: mild; 2: moderate; 3: severe; and 4: proliferative DR) to indicate different severity levels. The number of images in these classes are respectively 1805, 370, 999, 193, and 295. In the binary setting, class 0 is defined as the normal group, and the remaining classes (1-4) are defines as DR group which gives a split 1805:1857. The total number of training and test samples in the dataset were 3662 and 1928 respectively. However, the labels for the test samples were not released after the competition, so like other references, we used the available 3662 fundus images with labels. To ensure fair comparison with existing deep learning approaches, we employ a standard majority-based train–test split protocol. The class-wise distribution of samples in the training and test sets—before and after augmentation—is summarized in Table 1 for the binary classification setup and in Table 2 for the five-class (multiclass) setting.
@@ 4.2 Hyperparameters
All models were trained using the Adam optimizer with a fixed learning rate of $1\times 10^{-4}$ and a batch size of 32. Input images were resized to $\times 224$ 224 pixels and normalized using ImageNet statistics. Cross-entropy loss was used for both binary and multiclass classification tasks. Training was performed for a maximum of 500 epochs, with early stopping applied if the macro F1-score on the validation set did not improve for 50 consecutive epochs. During augmentation, each class in the training set was expanded to 20,000 samples to address class imbalance.
@@ 9 Future Work
Faisal Ahmed downloaded the data, conceptualized the study, prepared the code, performed the data analysis and wrote the manuscript. He reviewed and approved the final version of the manuscript.
@@
The author received no financial support for the research, authorship, or publication of this work.
Acknowledgement
The authors utilized an online platform to check and correct grammatical errors and to improve sentence readability.